OpenStudy (anonymous):

Explain the connection between the factors of a polynomial, zeros of a polynomial function, and solutions of a polynomial function.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

well as @satellite73 stated in your last problem, if x-2 is a factor for example, then x = 2 is a root since x-2 = 0 x-2+2 = 0+2 x = 2 in general, if x-k is a factor of p(x), then p(k) = 0 basically stating that x = k is a root of p(x)

jimthompson5910 (jim_thompson5910):

a "solution" to a polynomial is the same as a root or zero

OpenStudy (anonymous):

So the connection would be is that it's always a zero?

jimthompson5910 (jim_thompson5910):

yeah if x-3 is a factor, then x = 3 is a zero or root

OpenStudy (anonymous):

I don't get it

jimthompson5910 (jim_thompson5910):

ok let's say we had `x^2 + 5x + 6 = 0` how would we solve this?

OpenStudy (anonymous):

Solve for x

jimthompson5910 (jim_thompson5910):

one way is to factor, right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

what does x^2 + 5x + 6 factor to

OpenStudy (anonymous):

-6 and 1?

jimthompson5910 (jim_thompson5910):

find two numbers that multiply to 6 and add to 5

OpenStudy (anonymous):

3 and 2

jimthompson5910 (jim_thompson5910):

-6 times 1 = -6 but we want +6 as a result

jimthompson5910 (jim_thompson5910):

3 times 2 = 6 3 plus 2 = 5 works

jimthompson5910 (jim_thompson5910):

so x^2 + 5x + 6 factors to (x+3)(x+2)

OpenStudy (anonymous):

Oh I get it now

jimthompson5910 (jim_thompson5910):

if (x+3)(x+2) = 0 then either x+3 = 0 or x+2 = 0 that leads to x = -3 or x = -2

jimthompson5910 (jim_thompson5910):

what @satellite73 was doing in the last problem was thinking backwards from this process to construct a polynomial with the given roots

OpenStudy (anonymous):

Oh alright. So the connection between the three things is what?

jimthompson5910 (jim_thompson5910):

well in that example, x+3 is a factor so x=-3 is a root

jimthompson5910 (jim_thompson5910):

and x = -3 is a solution to the polynomial

OpenStudy (anonymous):

But I don't get how that's a connection to factors of polynomials, zeros of polynomials, and solutions of polynomial equation

jimthompson5910 (jim_thompson5910):

if you know a zero, then you can find the factor or vice versa

OpenStudy (anonymous):

If you answered that question how would you answer it

jimthompson5910 (jim_thompson5910):

just like that. That's how I would answer

OpenStudy (anonymous):

But I have to explain which sucks haha

OpenStudy (anonymous):

So how would you explain it. @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm basically repeating myself at this point. The explanation is all there. I explained it using an example. You factor x^2 + 5x + 6 to get (x+3)(x+2). Then the factors are used to find the roots using the zero product property

jimthompson5910 (jim_thompson5910):

You can take this process in reverse if they said something like "the roots are -3 and -2, find the polynomial"